OpenStudy (anonymous):

what is the probability of flipping 3 coins and getting exactly 1 head?

OpenStudy (anonymous):

how is it so though?

OpenStudy (anonymous):

let say 3 coins A B and C. among the 3 coins any 1 can show head. for each coin, probability of getting head= 1/2, and probability of getting tail=1/2. therefore probability=P(head) x P(tail)xP(tail) = 1/2 x 1/2 x 1/2 = 1/8. lolz sorry it should be 1/8.

OpenStudy (anonymous):

And I think that the coins are all flipped at once without considering the orders of flipping the coins. therefore we don't need to multiply the probability by 3.

OpenStudy (anonymous):

my computer quiz said it was 3/8.

OpenStudy (anonymous):

then I think the question should be asking to flip the coins one by one rather than flipping the coins all at once. so in this case it could be A be head, or B be head, or C be head. So just take the probability I found multiply by 3 where u get 3(1/8) = 3/8